OpenStudy (anonymous):

The sum of the finite series 4, 6, 4, 6 is:

ganeshie8 (ganeshie8):

add them ?

ganeshie8 (ganeshie8):

4+6+4+6

OpenStudy (anonymous):

so it should be 20

OpenStudy (anonymous):

they keep telling me its wrong

ganeshie8 (ganeshie8):

oh is there som trick let me think

OpenStudy (anonymous):

i think its fine what about this one...im so outdated

ganeshie8 (ganeshie8):

:) 20 is correct. ask them why they think its wrong

OpenStudy (anonymous):

what about the other problem

OpenStudy (anonymous):

is it 1/16

OpenStudy (anonymous):

or is it infinite

ganeshie8 (ganeshie8):

\(\large \sum \limits_{1}^4 \frac{1}{n}\)

OpenStudy (anonymous):

yep thats the problem

ganeshie8 (ganeshie8):

\(\large \sum \limits_{1}^4 \frac{1}{n}\) \(\large \frac{1}{1}+\frac{1}{2}+\frac{1}{3}+\frac{1}{4}\)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

LCD is 12

ganeshie8 (ganeshie8):

good :) keep going

OpenStudy (anonymous):

so how would you go about this one...The sum of the first three terms of 2 - n is:

ganeshie8 (ganeshie8):

is that \(2-n\), or \(\large2^{-n}\)

ganeshie8 (ganeshie8):

?